# Exploring Holomorphic Retracts

### v6

@@ 2.5. Fixed point sets and Retracts
While the above theorem allows retracts to be non-linear and we shall indeed furnish several examples of non-product (balanced, pseudoconvex) domains possessing non-linear retracts through origin in the form of graphs, we recall that such a phenomenon is precluded by Vesentini’s theorem in the convex case. Extremality boundary assumptions as in Vesentini’s theorem are indeed necessary as we shall see while establishing a converse to his theorem later. However, the hypothesis that the entire boundary be holomorphically extreme as in the extension of Vesetini’s theorem given by proposition 1.4 in [JJ] is inconvenient for discussing interesting intermediate examples in the sequel. This is addressed by proposition 1.4 which we now establish. Though this is in essence an imitation of the proof of the aforementioned proposition in [JJ], we prefer to include some details as this is a result about retracts that we shall need later in this article wherein the aforementioned relaxation will actually be used. And, also because it provides an answer towards the natural query as to whether an extension of Vesentini’s theorem holds if we add only an extremality hypothesis just as much as is needed about the boundary, but none about convexity in the setting of the above theorem 3.9.
@@
Proof of proposition 1.4: Choose any holomorphic retraction map $\rho:D\to D$ with $\rho(D)=Z$. We may observe by the above theorem 3.9, that the restriction $\rho|_{D_{L}}$ is a holomorphic map on $D_{L}$ whose image is precisely equal to $Z$; but we shall not need this for the rest of this proof here. Recall that $Z$ is a connected complex submanifold and the tangent space to $Z$ at the origin, is denoted by $L$; so, $L=T_{0}Z=\{v\in\mathbb{C}^{N}:\rho^{\prime}|_{D_{L}}(0)v=v\}$. Let $v\in D_{L},v\neq 0$. By our assumption on $\partial D_{L}$, it follows that the mapping
@@
$K_{D}\left(\rho|_{D_{L}}\circ\phi(0);(\rho|_{D_{L}}\circ\phi)^{\prime}(0)\right)=K_{D}\left(0,v/h_{D_{L}}(v)\right)=K_{D}\left(\phi(0);\phi^{\prime}(0)\right).$
Hence $\rho|_{D_{L}}\circ\phi$ is a $K_{D}$-geodesic for $(0,v/h_{D_{L}}(v))$. Because of our hypothesis, we have that $v/h_{D_{L}}(v)$ is a holomorphic extreme point of $D$. By uniqueness of complex geodesics in $D$ (exposited as proposition 11.3.5 in [Jrncki_invrnt_dst]), it follows that $\rho|_{D_{L}}\circ\phi=\phi\circ\psi$, where $\psi\in{\rm Aut}(\Delta)$. Substituting $\psi^{-1}(h_{D_{L}}(v))$ which is a point in $\Delta$, into the above equation, we see that $v$ lies in the image of $\rho|_{D_{L}}$. Hence $D_{L}\subset Z$. Now $T_{0}Z$ and $Z$ have the same dimension. Since $Z$ is connected, it follows from the principle of analytic continuation that $Z=D_{L}$. Hence $Z$ is linear. ∎
Some remarks about the foregoing pair of results are in order.
@@
To indicate just a few concrete examples to illustrate where proposition 1.4 may be applied (but not the other results in this subject directly at-least), we first deal with domains where the holomorphic extremality condition is neither violated maximally nor minimally; examples for ‘maximal’ violation is also considered but in section 8. Among the simplest of domains satisfying the hypotheses of proposition 1.4 in this intermediate fashion, we consider the domain $\Omega_{1}:=\Delta^{2}\cap D_{q}$, where $D_{q}:=\{(z,w)\in\mathbb{C}^{2}:|z|^{q}+|w|^{q}<2^{q}\}$ for $10<q<$ 1 is a dilated copy of the $\ell_{q}$-ball. Note that $\Omega_{1}$ is a bounded balanced pseudoconvex domain and $\partial\Omega_{1}$ fails to be holomorphically extreme (in its entirety); though of-course, there exists an open piece of $\partial\Omega_{1}$, which is holomorphically extreme. A combined application of theorem 1.10, theorem 1.13 and corollary 2.33, proves that every retract passing through origin is either a linear subspace of $\Omega_{1}$ or a graph of a non-linear holomorphic function over some linear subspace of $\Omega_{1}$.
@@
:=$\mathcal{E}:=\{z\in\mathbb{C}^{N}\;:\;|z_{1}|^{p_{1}}+\ldots+|z_{N}|^{p_{N}}<1\},$
where the exponents $p_{j}$’s are taken to be bigger than one and distinct (we disallow $N=2$ for that case is exceptional owing to the fact that all non-trivial retracts of convex domains in this case are at-most of dimension one and any pair of one-dimensional retracts are biholomorphic to $\Delta$). Then, for every $1<k<N$, there exist a pair of linear retracts of $\mathcal{E}$ of the same dimension $k$ which are diffeomorphic but biholomorphically inequivalent. This is easily seen for instance, by projecting $\mathcal{E}$ onto the span of different subsets of $k$-many of the standard basis vectors of $\mathbb{C}^{N}$ following the foregoing concrete examples.
Remark 3.13.
We shall assume until further notice that $D$ is convex. The existence in abundance of one dimensional retracts of strictly convex domains (in $\mathbb{C}^{N}$) is well-known due to the work of Lempert. Another well-known work which obtains retracts of higher dimensions under certain conditions is [Ian_Grhm] of Graham; as noted therein, the condition in the above theorem that $D_{L}=D\cap L$ be a linear retract of $D$ which is equivalent to the existence of a norm one projection onto $L$ (denoted by $\pi_{L}$ in the proof), is always met if $L$ is one dimensional – this follows by the Hahn–Banach theorem which may be used to first get a linear functional $f$ on $(\mathbb{C}^{N},\mu)$ of norm one which takes the value $1$ at a boundary point of $D_{L}$ (one may view $f$ as mapping $D$ onto $D_{L}\subset L\simeq\mathbb{C}$); the desired projection is then obtained by taking the linear projection of $\mathbb{C}^{N}$ onto $L$ along $\ker(f)$. On the other hand (also noted therein), if there exists a norm one projection of the Banach space $(\mathbb{C}^{N},\mu)$ onto a every linear subspace, then $(\mathbb{C}^{N},\mu)$ is actually a Hilbert space i.e., $\mu$ is actually induced by an inner product owing to classical results mentioned in the introduction namely that of Kakutani in [Kakutni] for the real case and Bohnenblust in [Bhnblst] for the complex case. Thus it must be noted that not every linear subspace intersected with $D$ need be a retract of $D$. But then again, we shall show in the next theorem that the retracts through the origin are all linear.
@@
Since $D_{L}$ is a linear retract of $D$, there exists a linear projection map $P$ such that $image(P)=D_{L}$. By idempotency of $P$, we can write $\mathbb{C}^{N}=L\oplus M$, where $M=\ker(P)$ and $L={\rm image}(I-P)$. Consider the map $\psi:D\to D$ defined by $\psi(z)=(T(z),\varphi(T(z)))$; this can be verified to be idempotent by a direct computation, thereby showing that $\psi$ is a holomorphic retraction from $D$ onto ${\rm Graph}(\varphi)$ over $D_{L}$, finishing the proof. ∎
It must be noted that the above does not mean that non-linear retracts automatically come into existence by taking the graph of any holomorphic function over any given linear retract of $D$ for, the graph can fail to be contained within $D$; that this is the only obstruction to generating non-linear retracts by graphs is the content of the above proposition. However, the obstruction is substantial: for instance, even in the simple case of $\mathbb{B}$, the standard unit ball with respect to the usual $l^{2}$-norm on $\mathbb{C}^{N}$, every retract through the origin is just a linear subspace and in fact, all retracts of $\mathbb{B}$ are affine-linear, as was first established by Rudin who then exposited it in chapter $8$ of his book [Rud_book_Fn_theory_unitball] as well. The proof of this, atleast as in [Rud_book_Fn_theory_unitball], uses the special feature of the $l^{2}$-norm that it comes from an inner product; however, this fact about the linearity of retracts through the origin is actually true for every bounded balanced convex domain with $\mathbb{C}$-extremal boundary. This is of-course the result of Vesentini which we extend and give a more elementary proof as already mentioned in the introduction i.e., we now lay down the proof of theorem 1.3.
Proof of theorem 1.3: Let $\rho:D\longrightarrow D$ be a retraction and $\rho(0)=0$. By lemma 3.7, $D\rho|_{0}$ is a projection of norm $1$ from $D$ onto $D_{L}=D\cap L$, where $L=T_{0}Z$. Note that $(D\rho|_{0})_{|_{D_{L}}}:D_{L}\longrightarrow D_{L}$ is the identity map on $D_{L}$ and that $D_{L}$ is a bounded balanced convex domain in $\mathbb{C}^{N}$. It follows from the proof of Krein-Milman theorem (the reader is referred to the book [Meggn] for the standard proof we have in mind here), that $D_{L}$ must have atleast one $\mathbb{R}$-extremal boundary point, say $z_{0}$. As $D\rho(0)$ acts as the identity map on $D_{L}$ (thereby in particular $D\rho(0)(z_{0})=z_{0}$), it follows that there is atleast one $\mathbb{R}$-extremal point in $D\rho(0)(D_{L})$ namely $z_{0}$ itself! To be precise $z_{0}$ is an $\mathbb{R}$-extremal point of $D$ not just $D_{L}$. Applying theorem 3.5 to the map $\rho_{|_{D_{L}}}$ viewed as a map from $D_{L}$ into the (larger) domain $D$, finishes the proof that $Z$ is linear.
@@
[…]
@@
$F_{i}(z)=\begin{cases}z_{i},\quad\qquad\qquad\qquad\qquad\text{if}~{}i\in J\\ c_{i,j_{1}}z_{j_{1}}+\ldots+c_{i,j_{r}}z_{j_{r}},\quad\text{if}~{}i\notin J,\\ \end{cases}$
@@
$w_{k}:=\begin{cases}\frac{|c_{m,j_{r}}|}{c_{m,j_{r}}},\ \text{if}\ k=j_{r}~{}\text{for some}~{}r\ \text{and}\ c_{m,j_{r}}\neq 0\\ 0,\quad\quad\text{otherwise}\end{cases}.$
@@
$F_{i}(e_{j_{k}})=\begin{cases}\delta_{i,j_{k}},\quad\qquad\text{if}~{}i\in J\\ c_{i,j_{k}},\qquad\text{if}~{}i\notin J,\\ \end{cases}$
@@
$F_{i}(z)=\begin{cases}z_{i},\quad\qquad\qquad\qquad\qquad\text{if}~{}i\in J\\ c_{i,j_{1}}z_{j_{1}}+\ldots+c_{i,j_{r}}z_{j_{r}},\quad\text{if}~{}i\notin J,\\ \end{cases}$
@@
As noted above, there exists a linear retract of $\Delta^{N}$ which is not spanned by mutually disjoint supported vectors. The converse implication however is true and can now be derived as a corollary.
@@
Without loss of generality assume that $\|v_{i}\|_{\infty}=1$ for $i=1,\dots,r$. There exists $J=\{j_{1},\ldots,j_{r}\}$ such that $\pi_{j_{i}}(v_{i})=1$, where $\pi_{i}$ denotes the projection to the $i^{th}$ component. We can write $v_{k}=e_{j_{k}}+w_{k}$. Note that $w_{k}$’s are disjointly supported vectors. If $m\in\{1,2,\ldots,N\}\setminus J$, then we define $c_{m,{j_{k}}}:=\pi_{m}(w_{k})$. Since $\text{supp}(w_{i})$’s are disjoint, we get $\pi_{m}(w_{k})\pi_{m}(w_{l})=0$ for $k\neq l$. Hence $\sum_{j\in J}|c_{m,j}|\leq 1$. By theorem 1.8, $Y=span\{v_{1},\ldots,v_{k}\}$ is a linear retract of $\Delta^{N}$. ∎
Turning now to the proof of theorem 1.15 and proceeding forth with it, as already remarked, the boundary of our product domain $D=B_{1}\times B_{2}$ does not enjoy the $\mathbb{R}$-extremality property. Nevertheless, we shall show that $Z$ is linear (indeed, $\rho(D_{L})=D_{L}=Z$) in several cases to be dealt with in the sequel and in all cases can be realized as holomorphic graphs over some subspace contained within the coordinate spaces $\mathbb{C}^{N_{1}}$ or $\mathbb{C}^{N_{2}}$ intercepted by $B_{1}$ or $B_{2}$ respectively. That is to say, while we already know that $\rho(D_{L})=Z$ is posited in $D$ as a holomorphic graph over $D_{L}$, what we shall be more precisely showing is that $Z$ can always be realized a holomorphic graph over some retract of $B_{1}$ or $B_{2}$ which in view of the last theorem is a linear subspace of the respective ball ($B_{1}$ or $B_{2}$). Indeed, we have finished our general observations and we shall now progress into the analysis of our specific $D$. First, a few words about notations about retracts $\rho$ of our specific product domain $D\subset\mathbb{C}^{N_{1}+N_{2}}$. When written out explicitly, $\rho$ has $N_{1}+N_{2}$ components: $(\rho_{1},\rho_{2},\ldots,\rho_{N_{1}},\rho_{N_{1}+1},\ldots,\rho_{N_{1}+N_{2}})$ where each of the $\rho_{j}^{\prime}s$ for $j=1,\ldots,N_{1}+N_{2}$ can apriori depend on all of the variables $w_{1},\ldots,w_{N_{1}},z_{1},\ldots,z_{N_{2}}$; so, we use $w=(w_{1},\ldots,w_{N_{1}})$ as our notation for the coordinates on $\mathbb{C}^{N_{1}}\supset B_{1}$ and $z=(z_{1},\ldots,z_{N_{2}})$ for that on $\mathbb{C}^{N_{2}}\supset B_{2}$.
@@
So $g_{1},g_{2}$ are holomorphic maps from the unit disc into the balls $B_{1},B_{2}$ respectively, mapping $0\in\Delta$ to the origin in the respective balls. We are thus essentially in the setting of the Schwarz lemma of Mazet recalled above as theorem 3.4, provided only that we verify that the derivatives of these maps at the origin satisfy the norm condition as in theorem 3.4. To do this, let us begin with the map $g_{1}$ whose derivative we compute using the chain-rule as:
@@
Then, using the facts that $d\pi_{1}\equiv\pi_{1}$ (as the projection $\pi_{1}$ is just linear), $\varphi$ maps $0\in\Delta$ to the origin in $B_{1}$ which inturn is fixed by $\rho$, and that $d\rho|_{0}$ is just the identity map on $T_{0}Z$ and that $(w^{0},z^{0})=d\varphi(0)\in T_{0}Z$, we get $dg_{1}(0)=w^{0}\in\partial B_{1}$, verifying the norm condition $\mu(w^{0})=1$ (and the $\mathbb{R}$-extremality of $dg_{1}(0)$, which is now trivial here owing to the hyporthesis that all boundary points of $B_{1}$ are $\mathbb{R}$-extremal). Hence, the aforementioned Schwarz lemma of Mazet, applied here to $g_{1}$ shows that that $\pi_{1}\circ\rho(tw_{0},tz^{0})=tw_{0}$, which means that $\rho^{1}:=\pi_{1}\circ\rho$ acts like the projection onto the $w$-axes on $D_{L}$:
@@
$w\longmapsto\big{(}\beta_{1}(w),\ldots,\beta_{N_{2}}(w)\big{)}$
is an isometry from the Banach space ($\pi_{1}(L)$, $\mu_{1}$) into the Banach space ($\mathbb{C}^{N_{2}}$, $\mu_{2}$) and therefore maps $B_{1}^{L}:=\pi_{1}(D_{L})$ which is the unit ball in $\pi_{1}(L)$ with respect to $\mu_{1}$ isometrically into $B_{2}$; the mapping is surjective if $N_{1}=N_{2}$ and $\pi_{1}(L)=\mathbb{C}^{N}$ where $N=N_{1}=N_{2}$. Hence, we see that this places strong restrictions about when this case-(I) can happen. For instance, this case cannot occur if $N_{1}=N_{2}=N>1$ and $B_{1}=\Delta^{N}$ is the standard unit polydisc and $B_{2}=\mathbb{B}$, the standard ball with respect to the $l^{2}$-norm in $\mathbb{C}^{N}$ and ${\rm dim}(D_{L})=N$; or more generally, case-(I) cannot arise as soon as $\pi_{1}(D_{L})$ is not biholomorphic to any of the linear sections of $B_{2}$. But then again, one must not be amiss to note that case-(I) does happen atleast once regardless of what particular domains $B_{1},B_{2}$ are, namely when ${\rm dim}(D_{L})=1$; indeed, the complex line segment spanned by every point in $\partial B_{1}\times\partial B_{2}$ is a retract (as is more generally true of the $\mathbb{C}$-line segment spanned by any boundary point).

### v7

@@ 2.9. Analytic and Semi-analytic sets
While the above theorem allows retracts to be non-linear and we shall indeed furnish several examples of non-product (balanced, pseudoconvex) domains possessing non-linear retracts through origin in the form of graphs, we recall that such a phenomenon is precluded by Vesentini’s theorem in the convex case when the boundary is $\mathbb{C}$-extremal. That such extremal boundary conditions are indeed necessary will be shown in course of establishing a converse to Vesentini’s theorem later. However, the hypothesis that the entire boundary be holomorphically extreme as in the extension of Vesentini’s theorem given by proposition 1.4 in [JJ] is inconvenient for discussing interesting intermediate examples in the sequel. This is addressed by proposition 1.3 whose proof we now lay down. Though this can be done by an imitation of the proof of the aforementioned proposition in [JJ], we prefer to include some details as this is a result about retracts and further, an occassion will arise later in this article wherein the aforementioned relaxation will actually be used. It also provides an answer towards the natural query as to whether an extension of Vesentini’s theorem holds if we add only an extremality hypothesis just as much as is needed about the boundary, but none about convexity in the setting of the above theorem 3.10.
@@
Proof of proposition 1.3: This is one of the few occasions in this article wherein we discuss $2$ proofs, with a one sentence proof following this, to show a possible advantage of circumventing geodesics at times, as mentioned in the introduction. Choose any holomorphic retraction map $\rho:D\to D$ with $\rho(D)=Z$. We may observe by the above theorem 3.10, that the restriction $\rho_{|_{D_{L}}}$ is a holomorphic map on $D_{L}$ whose image is precisely equal to $Z$; but we shall not need this for the rest of this proof here. Recall that $Z$ is a connected complex submanifold and the tangent space to $Z$ at the origin, is denoted by $L$; so, $L=T_{0}Z=\{v\in\mathbb{C}^{N}:\rho^{\prime}_{|_{D_{L}}}(0)v=v\}$. Let $v\in D_{L},v\neq 0$. By our assumption on $\partial D_{L}$, it follows that the mapping
@@
$K_{D}\left(\rho_{|_{D_{L}}}\circ\phi(0);(\rho_{|_{D_{L}}}\circ\phi)^{\prime}(0)\right)=K_{D}\left(0,v/h_{D_{L}}(v)\right)=K_{D}\left(\phi(0);\phi^{\prime}(0)\right).$
Hence $\rho_{|_{D_{L}}}\circ\phi$ is a $K_{D}$-geodesic for $(0,v/h_{D_{L}}(v))$. Because of our hypothesis, we have that $v/h_{D_{L}}(v)$ is a holomorphic extreme point of $D$. By uniqueness of complex geodesics in $D$ (exposited as proposition 11.3.5 in [Jrncki_invrnt_dst]), it follows that $\rho_{|_{D_{L}}}\circ\phi=\phi\circ\psi$, where $\psi\in{\rm Aut}(\Delta)$. Substituting $\psi^{-1}(h_{D_{L}}(v))$ which is a point in $\Delta$, into the above equation, we see that $v$ lies in the image of $\rho_{|_{D_{L}}}$. Hence $D_{L}\subset Z$. Now $T_{0}Z$ and $Z$ have the same dimension. Since $Z$ is connected, it follows from the principle of analytic continuation that $Z=D_{L}$. Hence $Z$ is linear. ∎
An alternative short proof of this proposition is obtained simply by noting that it is a direct corollary of theorem 3.6; specifically, by applying the second part of that theorem to any one of the holomorphic retraction maps $\rho$ restricted to $D_{L}$. ∎
Some remarks about the foregoing pair of results are in order. Firstly, a remark about the last proof: note that the arguments up-till the conclusion that ${\rm image}(\tilde{\varphi})\subset\partial D$, goes through in the setting of above theorem (same as theorem 1.15), whose statement we may therefore refine as follows. If we call the largest complex analytic variety within $\partial D$ containing a point $p\in\partial D$ as the ‘holomorphic face’ of $\partial D$ at $p$, then we may say that the mapping (denote it by $F$ say), whose graph is the retract $Z$, as in theorem 1.15, takes values within the union of such holomorphic faces at boundary points $p$, for $p$ varying in $\partial D_{L}=\partial D\cap L$. As we already know that the range of the map $F$ is contained in $\ker(D\rho(0))$, the above helps refine this information about ${\rm range}(F)$. We thereby get another proof of proposition 1.3 as an immediate corollary of the foregoing theorem. This owes to the assumption that the holomorphic face on $\partial D_{L}$ becoming a singleton in the situation of that proposition.
@@
To indicate just a few concrete examples to illustrate where proposition 1.3 may be applied (but not the other results in this context directly at-least), we first deal with domains where the holomorphic extremality condition is neither violated maximally nor minimally; examples for ‘maximal’ violation is also considered but in section 8. Among the simplest of domains satisfying the hypotheses of proposition 1.3 in this intermediate fashion, we consider the domain $\Omega_{1}:=\Delta^{2}\cap D_{q}$, where $D_{q}:=\{(z,w)\in\mathbb{C}^{2}:|z|^{q}+|w|^{q}<2^{q}\}$ for $0<q<1$ is a dilated copy of the $\ell_{q}$-ball. Note that $\Omega_{1}$ is a bounded balanced pseudoconvex domain and $\partial\Omega_{1}$ fails to be holomorphically extreme; though of-course, there exists an open piece of $\partial\Omega_{1}$, which is holomorphically extreme. A combined application of theorem 1.12, theorem 1.15 and corollary 2.38, proves that every retract passing through origin is either a linear subspace of $\Omega_{1}$ or a graph of a non-linear holomorphic function over some linear subspace of $\Omega_{1}$.
@@
$\mathcal{E}:=\{z\in\mathbb{C}^{N}\;:\;|z_{1}|^{p_{1}}+\ldots+|z_{N}|^{p_{N}}<1\},$
where the exponents $p_{j}$’s are taken to be bigger than one and distinct (we disallow $N=2$ for that case is exceptional owing to the fact that all non-trivial retracts of convex domains in this case are at-most of dimension one and any pair of one-dimensional retracts are biholomorphic to $\Delta$). Then, for every $1<k<N$, there exists a pair of linear retracts of $\mathcal{E}$ of the same dimension $k$ which are diffeomorphic but biholomorphically inequivalent. This is easily seen for instance, by projecting $\mathcal{E}$ onto the span of different subsets of $k$-many of the standard basis vectors of $\mathbb{C}^{N}$ following the foregoing concrete examples.
@@
Since $D_{L}$ is a linear retract of $D$, there exists a linear projection map $P$ such that $P(D)=D_{L}$. By idempotency of $P$, we can write $\mathbb{C}^{N}=L\oplus M$, where $M=\ker(P)$ and $L={\rm image}(P)$. Consider the map $\psi:D\to D$ defined by $\psi(z)=(P(z),\varphi(P(z)))$; this can be verified to be idempotent by a direct computation, thereby showing that $\psi$ is a holomorphic retraction from $D$ onto ${\rm Graph}(\varphi)$ over $D_{L}$, finishing the proof. ∎
It must be noted that the above does not mean that non-linear retracts automatically come into existence by taking the graph of any holomorphic function over any given linear retract of $D$ for, the graph can fail to be contained within $D$; that this is the only obstruction to generating non-linear retracts by graphs is the content of the above proposition. However, the obstruction is substantial: for instance, even in the simple case of $\mathbb{B}$, the standard unit ball with respect to the usual $l^{2}$-norm on $\mathbb{C}^{N}$, every retract through the origin is just a linear subspace and in fact, all retracts of $\mathbb{B}$ are affine-linear, as was first established by Rudin who then exposited it in chapter $8$ of his book [Rud_book_Fn_theory_unitball] as well. The proof of this, at least as in [Rud_book_Fn_theory_unitball], uses the special feature of the $l^{2}$-norm that it comes from an inner product; however, this fact about the linearity of retracts through the origin is actually true in much greater generality as stated in proposition 1.3 .
@@
Proof of proposition 1.7: Assume $L$ is a linear retract of $\Delta^{N}$. We can write $L=Y\cap\Delta^{N}$, for some linear subspace $Y\subset\mathbb{C}^{N}$. By theorem-3 in [Heath_Suff], there exists $J=\{j_{1},j_{2},\ldots,j_{r}\}$ and holomorphic functions $F_{i}:M\to\Delta,~\ 1\leq i\leq N$ where $M=\{z\in\Delta^{N}:z_{j}=0~\text{if}~j\notin J\}$ such that $F_{i}(z)=z_{i}$ if $i\in J$ and
@@
[…]
@@
$w_{k}:=\begin{cases}\frac{|c_{m,j_{i}}|}{c_{m,j_{i}}},\ \text{if}\ k=j_{i}~\text{for some}~i\ \text{and}\ c_{m,j_{i}}\neq 0\\ 0,\quad\quad\text{otherwise}\end{cases}.$
@@
[…]
@@
[…]
@@
As noted above, there exists a linear retract of $\Delta^{N}$ which is not spanned by mutually disjoint supported vectors. The converse implication however, is true and can now be derived as a corollary.
@@
Without loss of generality assume that $\|v_{i}\|_{\infty}=1$ for $i=1,\dots,r$. There exists $J=\{j_{1},\ldots,j_{r}\}$ such that $|\pi_{j_{i}}(v_{i})|=1$, where $\pi_{i}$ denotes the projection to the $i^{th}$ component. After multiplying an unimodular constant by $v_{i}$, we may assume that $\pi_{j_{i}}(v_{i})=1$. So, we can write $v_{k}=e_{j_{k}}+w_{k}$. Note that $w_{k}$’s are disjointly supported vectors. If $m\in\{1,2,\ldots,N\}\setminus J$, then we define $c_{m,{j_{k}}}:=\pi_{m}(w_{k})$. Since $\text{supp}(w_{i})$’s are disjoint, we get that $\pi_{m}(w_{k})\pi_{m}(w_{l})=0$ for $k\neq l$. Hence $\sum_{j\in J}|c_{m,j}|\leq 1$. By proposition 1.7, $Y=span\{v_{1},\ldots,v_{k}\}$ is a linear retract of $\Delta^{N}$. ∎
Proposition 3.16.
Let $W$ be a linear subspace of $\mathbb{C}^{N}$. Then $W\cap\Delta^{N}$ is a $k$-dimensional linear retract of $\Delta^{N}$ if and only if $W$ is isometric to $(\mathbb{C}^{k},\|\cdot\|_{\infty})$.
Proof.
One part of the implication will need tools that will be discussed only later below. Namely: if $W$ is isometric to $(\mathbb{C}^{k},\|\cdot\|_{\infty})$, then by proposition 5.4 below, $W\cap\Delta^{N}$ is a $k$-dimensional linear retract of $\Delta^{N}$. We shall therefore now deal with the converse here; to do this, assume that $W$ is a $k$-dimensional linear retract of $\Delta^{N}$. By proposition 1.7, there exists $J=\{j_{1},\ldots,j_{k}\}\subset\{1,2,\ldots,N\}$ such that $L$ is spanned by $v_{1},\ldots,v_{k}$, where each $v_{r}$ is of the form
$v_{r}=\left(\sum_{i\notin J}c_{i,j_{r}}e_{i}\right)+e_{j_{r}}$
and for each $m\in\{1,2,\ldots,N\}\setminus J$ satisfies $\sum_{j\in J}|c_{m,j}|\leq 1$. Without loss of generality assume that $J=\{j_{1},\ldots,j_{k}\}=\{1,2,\ldots,k\}$. Consider the linear map $S:W\to\mathbb{C}^{k}$ defined by $S(v_{i})=e_{i}$ for $i=1,2,\ldots,k$ and extended to other vectors of $W$ by linearity. Now we claim that $S$ is an isometry. To prove this claim, write each vector $v\in W$ as $v=\lambda_{1}v_{1}+\ldots+\lambda_{k}v_{k}$ for some $\lambda_{i}\in\mathbb{C}$. Note that $S(v)=(\lambda_{1},\ldots,\lambda_{k})$. Choose $i_{0}\in\{1,2,\ldots,k\}$ such that $\|(\lambda_{1},\ldots,\lambda_{k})\|_{\infty}=|\lambda_{i_{0}}|$. Hence $\|S(v)\|=|\lambda_{i_{0}}|$. Using the expression of $v_{r}$, we can expand $v$ as
$v=\sum_{l=1}^{k}\sum_{i=k+1}^{N}\lambda_{l}(c_{i,l}e_{i}+e_{l})$
Observe that the first $k$ components of $v$ are $\lambda_{1},\ldots,\lambda_{k}$ and for each $i=k+1~\text{to}~N$ , $i^{\text{th}}$ component of $v$ is $\sum_{l=1}^{k}\lambda_{l}c_{i,l}$. Note that for each $i=k+1$ to $N$, we have
$\left|\sum_{l=1}^{k}\lambda_{l}c_{i,l}\right|\leq|\lambda_{i_{0}}|\left(\sum_{l=1}^{k}|c_{i,l}|\right)\leq|\lambda_{i_{0}}|$
Hence we obtain that $\|S(v)\|_{\infty}=\|v\|_{\infty}$. This finishes the proof that $S$ is an isometry. ∎
We turn now to the proof of theorem 1.17. First, a few words about notations about retractions $\rho$ specific to this theorem. When written out explicitly, $\rho$ has $N_{1}+N_{2}$ components: $(\rho_{1},\rho_{2},\ldots,\rho_{N_{1}},\rho_{N_{1}+1},\ldots,\rho_{N_{1}+N_{2}})$ where each of the $\rho_{j}^{\prime}s$ for $j=1,\ldots,N_{1}+N_{2}$ can apriori depend on all of the variables $w_{1},\ldots,w_{N_{1}},z_{1},\ldots,z_{N_{2}}$; so, we use $w=(w_{1},\ldots,w_{N_{1}})$ as our notation for the coordinates on $\mathbb{C}^{N_{1}}\supset B_{1}$ and $z=(z_{1},\ldots,z_{N_{2}})$ for that on $\mathbb{C}^{N_{2}}\supset B_{2}$. We denote the Minkowski functionals of $B_{1},B_{2}$ by $\mu_{1},\mu_{2}$ respectively. We spend some words for the relevant convex case i.e., the special case when $\mu_{1},\mu_{2}$ norms such that their unit balls are strictly convex – we refer to this as ‘the convex case’ for brevity, till the proof of theorem 1.17 is completed. This explicit spelling out is for the convenience of future use of this result in conjunction with results on norm one projections in Banach spaces from the functional analysis literature.
@@
So $g_{1},g_{2}$ are holomorphic maps from the unit disc into the balls $B_{1},B_{2}$ respectively, mapping $0\in\Delta$ to the origin in the respective balls. We are thus essentially in the setting of theorem 3.6 above (in-case $B_{1},B_{2}$ are strictly convex, the Schwarz lemma of Mazet recalled above as theorem 3.4 may also be applied), provided only that we verify that the derivatives of these maps at the origin satisfy the extremality condition as required. To do this, let us begin with the map $g_{1}$ whose derivative we compute using the chain-rule as:
@@
Then, using the facts that $d\pi_{1}\equiv\pi_{1}$ (as the projection $\pi_{1}$ is just linear), $\varphi$ maps $0\in\Delta$ to the origin in $B_{1}$ which inturn is fixed by $\rho$, and that $d\rho|_{0}$ is just the identity map on $T_{0}Z$ and that $(w^{0},z^{0})=d\varphi(0)\in T_{0}Z$, we get $dg_{1}(0)=w^{0}\in\partial B_{1}$ which is a holomorphically extreme boundary point by hypothesis. The aforementioned theorem applied to $g_{1}$ shows that that $\pi_{1}\circ\rho(tw_{0},tz^{0})=tw_{0}$, which means that $\rho^{1}:=\pi_{1}\circ\rho$ acts like the projection onto the $w$-axes on $D_{L}$:
@@
(3.4) $w\longmapsto\left(\beta_{1}(w),\ldots,\beta_{N_{2}}(w)\right)$
maps $B_{1}^{L}:=\pi_{1}(D_{L})$ into $B_{2}$.
In this paragraph, we remark about the special case when $B_{1},B_{2}$ are both strictly convex. We may phrase the last observation as saying that the map (3.4) is an isometry from the Banach space ($\pi_{1}(L)$, $\mu_{1}$) into the Banach space ($\mathbb{C}^{N_{2}}$, $\mu_{2}$) and therefore maps $B_{1}^{L}$ which is the unit ball in $\pi_{1}(L)$ with respect to the norm $\mu_{1}$ isometrically into $B_{2}$; the mapping is surjective if $N_{1}=N_{2}$ and $\pi_{1}(L)=\mathbb{C}^{N}$ where $N=N_{1}=N_{2}$. Hence, we see that this places strong restrictions about when this case-(I) can happen. For instance, this case cannot occur if $N_{1}=N_{2}=N>1$ and $B_{1}=\Delta^{N}$ is the standard unit polydisc and $B_{2}=\mathbb{B}$, the standard ball with respect to the $l^{2}$-norm in $\mathbb{C}^{N}$ and ${\rm dim}(D_{L})=N$; or more generally, case-(I) cannot arise as soon as $\pi_{1}(D_{L})$ is not biholomorphic to any of the linear sections of $B_{2}$. But then again, one must not be amiss to note that case-(I) does happen at least once regardless of what particular balls $B_{1},B_{2}$ are, namely when ${\rm dim}(D_{L})=1$; indeed, the complex line segment spanned by every point in $\partial B_{1}\times\partial B_{2}$ is a retract (as is more generally true of the $\mathbb{C}$-line segment spanned by any boundary point of $D$, as we are dealing only with the convex case in this para).